OpenStudy (anonymous):

Use substitution to solve each system of equations. x=7y-4 2x-3y=14

hero (hero):

Hint: Plug the expression for x into the second equation, then solve for y: 2x - 3y = 14 2(7y - 4) - 3y = 14

OpenStudy (anonymous):

would it go to 2x-7y=18?

hero (hero):

Where'd you get that bro? I'm not seeing how you got that from the original equations.

OpenStudy (anonymous):

hold on i messed up